Suppose M is between L and N. Use the segment addition postulate to solve for the variable. Then find the indicated lengths.

LM= 5x+9, MN= 2x-6, LN = 59

find x, LM, and MN please. Please show work. this is to study for a big test tomorrow.

Answer by stanbon(75887)

Suppose M is between L and N. Use the segment addition postulate to solve for the variable. Then find the indicated lengths.
LM= 5x+9, MN= 2x-6, LN = 59
---
Draw the line segment LN. Put M somewhere between L and N.
Label LM as 5x+9 ; Label MN as 2x-6
Label the length of LN as 59
-----------------------------------------
find x, LM, and MN please.
Equation:
LM + MN = LN
5x+9 + 2x-6 = 59
7x + 3 = 59
7x = 56
x = 8
------------------------------
LM = 5x+9 = 5*8+9 = 49
MN = 2x-6 = 2*8-6 = 10
==============================
Cheers,
Stan H.